# End-to-End 4D Heart Mesh Recovery Across Full-Stack and Sparse Cardiac MRI

### v1

@@ Abstract
Reconstructing cardiac motion from cine CMR sequences is critical for diagnosis, prediction, and intervention. Existing methods rely on complete CMR stacks to infer full heart motion, limiting their utility in intra-procedural scenarios where only sparse observations are available. We present TetHeart, the first end-to-end framework that unifies full 4D multi-structure heart mesh recovery from both offline full-stack acquisitions and intra-procedural sparse-slice observations. Our method leverages deep deformable tetrahedra, an explicit–implicit hybrid representation, to capture shape and motion in a coherent space shared across cardiac structures. It is initialized from high-quality pre-procedural or offline-acquired full stacks to build detailed, patient-specific heart meshes, which can then be updated using whatever slices are available, from full stacks down to a single slice. We further incorporate several key innovations: (i) an attentive mechanism for slice-adaptive 2D–3D feature assembly that dynamically integrates information from arbitrary numbers of slices at any position, combined with a distillation strategy from full-slice to sparse-slice settings to ensure accurate reconstruction under extreme sparsity; and (ii) a two-stage weakly supervised motion learning scheme requiring only keyframe (e.g., ED and ES) annotations. Trained and validated on three large public datasets and externally evaluated zero-shot on additional private interventional and public CMR datasets, TetHeart achieves state-of-the-art accuracy and strong generalization in both pre- and intra-procedural settings.
{IEEEkeywords}
cardiac magnetic resonance (CMR), deformable tetrahedra, explicit–implicit hybrid representation, shape and motion modeling, sparse observation.
@@ 1 Introduction
The heart is one of the most vital organs in the human body [1], and accurate modeling of its shape and motion is central to numerous medical applications ranging from diagnosis to therapy planning [2, 3, 4]. Cardiac magnetic resonance (CMR) offers rich temporal imaging of cardiac anatomy and has become an indispensable modality for both shape [5, 6, 7, 8, 9] and motion [10, 11, 12, 13, 14] modeling. However, to this day, such modeling remains confined to offline analysis and pre-operative planning. Existing techniques [5, 6, 7, 8, 9, 10, 11, 12, 13, 14] cannot be applied during an intervention, as they require full-stack CMR data that cannot be acquired in real time. In the intervention room, only a handful of real-time slices can be captured, usually at lower resolution [15, 16, 17], a limitation also confirmed in routine practice at our collaborating university hospital. This scenario motivates us to move beyond the current state of the art and recover full 3D shape and motion from sparse and partial observations. This will enable real-time tracking of patient-specific heart dynamics, with a view to delivering guidance to interventional teams.
To this end, we introduce TetHeart, a unified framework for reconstructing the full 4D motion of multiple cardiac structures from either complete CMR stacks or intra-procedural sparse data. As in Fig. 1, TetHeart first leverages high-quality pre-operative or offline-acquired CMR stacks to build a detailed, patient-specific heart mesh, and then updates it in an end-to-end manner using any available slices, from a complete stack down to a single slice. To make this possible, TetHeart builds on deep deformable tetrahedra [18], an explicit–implicit hybrid representation that unites the optimization efficiency of signed distance functions with the geometric manipulability of meshes, forming the foundation for jointly modeling shape and motion in a coherent space shared across structures. On top of this, our first core contribution is an attentive mechanism for slice-adaptive 2D-3D feature assembly, which enables the model to handle sparse input by dynamically integrating information from arbitrary slices at arbitrary positions. Additionally, we introduce a distillation strategy from full-slice to sparse-slice settings, ensuring accurate reconstruction under extreme input sparsity. Our second contribution is a two-stage weakly supervised motion learning scheme that requires only keyframe (e.g., ED and ES) annotations, making it feasible to train TetHeart across diverse clinical environments. The core idea is to first train a static shape reconstruction branch on labeled frames, and then repurpose it to initialize shapes for unlabeled frames, from which motion dynamics are recovered on the labeled frames to provide supervision.
@@
Figure 1: TetHeart reconstructs cardiac motion from observed 2D slices. At $t=0$, a 3D CMR stack $\mathbf{I}^{0}$ is processed by the AFA module with self-attention to extract a volumetric static feature $V_{\text{static}}^{0}$, which is then used to generate the initial tetrahedra $\mathcal{G}^{0}$. For each subsequent time step $t$, the observed 2D slices $\mathbf{O}^{t}$ are encoded via the AFA module with cross-attention to obtain a volumetric motion feature $V_{\text{motion}}^{t}$, which is then used to recover motion by predicting offsets that deform $\mathcal{G}^{0}$. Weights are shared between the static reconstruction branch and dynamic deformation branch.
By representing the heart in a shared deformable space, TetHeart preserves spatial alignment between structures while adapting to varying sparsity levels. This offline/online capability enables comprehensive, patient-specific motion tracking across the entire clinical workflow: not only achieving state-of-the-art accuracy for offline use with full-stack CMR, but also, more importantly, unlocking online use in intra-procedural scenarios where only sparse observations are available.
We conducted extensive experiments to train and validate our model on three large public datasets: ACDC [19], M&Ms [20], and M&Ms-2 [21]. Remarkably, Remarkably, with a single set of weights learned from only end-diastolic (ED) and end-systolic (ES) annotations, our method generalizes seamlessly across both online few-slice and offline full-slice settings, consistently outperforming existing approaches not only on geometric metrics (Chamfer Distance and Dice) but also on clinically relevant indexes related to ejection fraction. We further performed external evaluations without retraining to assess generalizability. First, we tested the method on a privately collected dataset tailored for the online interventional MRI scenario at our collaborating cardiac CMR center, demonstrating its potential for deployment in real-world clinical environments. In addition, we evaluated the model on the public 4DM dataset [14], which provides complete temporal annotations of the myocardium, thereby demonstrating that the proposed weakly supervised motion learning scheme enables accurate reconstruction across the entire cardiac cycle. Together, these results highlight the effectiveness, robustness, and clinical promise of TetHeart. We will make our code public.
@@ 2.1 Static Cardiac Shape Modeling
Reconstructing the heart from 3D medical images holds significant clinical value. However, cardiac MR (CMR) images often exhibit low through-plane resolution, which makes accurate reconstruction a challenge. Over the years, many approaches to addressing it have been proposed [22, 5, 23, 24, 25, 26, 27, 28, 29, 9] . They can be categorized into traditional optimization-based techniques and deep learning-based methods.
2.1.1 Optimization-Based Approaches
They usually rely on a two-stage pipeline. Contours are first extracted from the CMR images—either manually or using a segmentation algorithm—and then a template mesh is optimized to fit these contours. For example, the approach of [23] starts from a predefined tubular initial mesh that is deformed to minimize a contour-matching loss. More recently, a 3D active shape model (3D-ASM) and an image intensity model was used in [29] to align initial meshes with ground-truth contours. While these approaches can produce good results, they also have drawbacks. As they rely on test-time optimization, the reconstruction process is often time-consuming, taking from tens of seconds to several minutes per mesh. When applied to a full image sequence, this may extend to over an hour, making them impractical for real-time applications. Moreover, these methods’ performance has only been evaluated using ground-truth contours, which require labor-intensive manual annotation. What their performance is when using predicted contours instead remains an open question.
2.1.2 Deep Learning-Based Approaches
Deep learning-based methods leverage prior knowledge learned from large datasets to accelerate mesh reconstruction. For instance, in [28], sets of contours are treated as sparse point cloud and decoded into 3D meshes using point- and graph-convolutions. Similarly, in [26], contour points are embedded into a 3D volume and a hybrid architecture made of 3D CNNs and GCNs is used to progressively deform a template mesh. These methods increase speed but also suffer from several limitations. Because meshes are reconstructed solely from extracted contours or points, the rich appearance information contained in cine images is not fully exploited. Furthermore, as contour extraction is not necessarily perfect, such two-states approaches can result in error accumulation [26].
@@ 2.2 Dynamic Cardiac Motion Modeling
Arguably, when reconstructing deforming hearts from sequences of CMR scans, one could simply run a static reconstruction algorithm on each individual frame. However, this would fail to exploit temporal consistency and make it difficult to establish inter-frame correspondences. Thus a number of motion-based approaches that compute the deformation from frame to frame have been proposed.
2.2.1 Motion from Full-Slice Offline MRI Sequences
@@
When full CMR stacks are available, recent deep learning-based approaches [11, 12, 13, 14, 6] can be used. MulViMotion [12] trains 3D CNNs to estimate left ventricular myocardial motion by predicting voxel-based deformation fields from CMR data, and then building 3D meshes by warping segmentation masks. DeepMesh [13] decouples mesh reconstruction from motion estimation, and uses 3D CNNs and bilinear interpolation to predict per-vertex deformation. 4DMR [14] models the myocardium using implicit representation. Decoupled shape and motion latent codes are optimized during test time to predict an initial shape and the backward motion fields. We use MulViMotion, DeepMesh and 4DMR as baselines because they are among the best currently available.
@@
Unfortunately, these approaches also suffer from several limitations. First, some of them focus only on myocardium, neglecting other cardiac structures, preventing comprehensive heart assessment. Second, optimization-based methods like 4DMR are time-consuming at inference time. And, finally and most importantly, they require full scans that cannot be acquired at sufficient frame rate, which precludes intra-procedural use as discussed below.
2.2.2 Motion from Sparse Online MRI Sequences
@@
CMR imaging has a reputation for being a ”slow” modality. Acquiring a full MRI stack involves sequential slice acquisition over a 4–6 heartbeat breath-hold, which can take several minutes [16, 17]. While this is feasible for pre-interventional planning, it is not suitable for intra-procedural settings.
@@
Recently, real-time MRI (RT-MRI) has gained favor due to its ability to capture dynamic processes without the need for gating, synchronization, or repetition [16]. Its applications include complex electrophysiology procedures [30], congenital heart disease [31], and other cardiac interventions [15, 32]. RT-MRI provides excellent soft-tissue contrast for assessing interventions like ablations [33, 17] and allows for continuous visualization and tracking of devices, and it is free of ionizing radiation [30, 34].
@@
Nevertheless, although RT-MRI enables rapid and continuous image acquisition, it still involves a fundamental tradeoff between spatial resolution, temporal resolution, artifacts and reconstruction latency [16]. Therefore, achieving a sufficient temporal resolution (e.g., 20–25 Hz) to be used during invasive procedures, can only be done at the cost of acquiring a limited number of 2D slices at lower spatial resolution [15, 32, 34, 16, 17]. We are not aware of any current method that can reconstruct 3D cardiac motion using this kind of data. Yet, delivering a high-quality, dynamic 3D heart model to the intervention room based on real-time observations is crucial for guiding high-precision minimally invasive cardiac procedures [35, 36, 15] and for real-time localization of devices within the heart [30, 37]. This is the application that motivated us to begin this study.
@@ 3 Method
In this section, we introduce TetHeart, our framework for accurate 4D mesh reconstruction from either full CMR scans or sparse sets of 2D slices, as depicted by Fig 1. We first introduce our heart presentation that is based on deep deformable tetrahedra [18], an explicit–implicit hybrid representation that unites the optimization efficiency of signed distance functions with the geometric manipulability of meshes. We then introduce our Attentive 2D–3D Feature Assembler (AFA) module, which enables effective feature encoding from 2D slices with varying numbers and locations. Finally, we present our weakly supervised training scheme that makes it possible to train given only a limited number of annotations.
@@ 3.1 Formalization
To model deforming hearts, we use the deep deformable tetrahedra [18] formalism. We discretize the 3D space using a deformable tetrahedral grid, in which each vertex possesses an SDF value, denoted as $\mathcal{G}=(\{\mathbf{v}_{i}\},\{s_{i}\},\mathcal{T})$, where $\mathbf{v}_{i}$ and $s_{i}$ are the vertices and their corresponding SDF values in $\mathcal{T}$, the set of all tetrahedrons. The explicit surface can then be extracted using the differentiable marching tetrahedra algorithm [38].
Given an initial 3D heart model $\mathcal{G}^{0}$ reconstructed from the end-diastolic (ED) frame $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$, our goal is to estimate a deformed model $\mathcal{G}^{t}$ at each time instant $t$, based on subsequent observations $\mathbf{O}^{t}$ that reflect the ongoing cardiac motion. We do this by updating the vertices of $\mathcal{G}^{0}$. We write
@@
    Online. For intra-procedural use, $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$ is a full-slice image acquired before the intervention, taking as much time as necessary. However, it has to be possible to acquire $\mathbf{O}^{t}$ at sufficient frame rate and, consequently, it can only comprise a small set—from 1 to 3—of 2D slices at lower spatial resolution and at arbitrary locations [15].
@@ 3.2 Attentive 2D-3D Feature Assembler
Because the observations can be either a full CMR stack, a sparse set of slices, or anything in between, we need a mechanism to extract useful features in all cases. To this end, we developed the Attentive 2D-3D Feature Assembler (AFA) module that relies on an attention mechanism that takes $\mathbf{I}^{0}$ and $\mathbf{O}^{t}$ as input and produces features.
Let us write $\mathbf{O}^{t}$ as $\{\mathbf{I}^{t}_{s}\in\mathbb{R}^{H\times W}|s=1,\ldots,S\}$, a collection of 2D slices where $S$ is the total number of slices. With this notation, $\mathbf{I}^{0}$ can be written as $\{\mathbf{I}^{0}_{d}\in\mathbb{R}^{H\times W}|d=1,\ldots,D\}$. A full-slice observation is simply one where $S=D$.
@@
Motion Features. A naive way to use an attention mechanism to construct motion features would be to treat the $F^{0}_{2d}$ as the query and the $F^{t}_{2d}$ as both keys and values. However, this would be inefficient because global attention aims at capturing long-range dependencies, which is not needed to model the inherently local nature of anatomical motion. Also, this requires computing full pairwise relationships between all query and key positions, which is computationally expensive when using many slices and high-resolution feature maps.
To avoid this, we take our inspiration from the Swin Transformer [39] and use instead a variant of attention that preserves locality and significantly reducing computational complexity. For each voxel in $F^{0}_{2d}$, we first identify $k_{1}$ spatially nearest 2D slices, or all slices if there are only $k_{1}$ or less. Within each selected slice, we then retrieve the features at the $k_{2}$ spatially closest positions to serve as keys and values. Attention is subsequently computed only between the query and its corresponding set of localized keys and values. This can be written as
$V^{t}_{\rm motion}=AFA(F^{0}_{2d},F^{t}_{2d})=\text{MHAttention}(Q,K,V)\;,$ (2)
@@
where NN-Select denotes the nearest slice and position selecting operation. MHAttention denotes a Multi-Head Attention [40] operation. A learnable position embedding is used to encode the spatial information into the calculation of attention. In practice, we set $k_{1}=3$ and $k_{2}=9$. This configuration ensures that the computational complexity of our attention mechanism is approximately equivalent to that of a 3D convolution with a kernel size of 3. As a result, the AFA module achieves high computational efficiency while retaining the flexibility and adaptability of attention-based feature aggregation.
While our AFA module effectively encodes features from varying 2D slice configurations, a persistent challenge remains in the online scenario, where the limited number of slices can hinder accurate 3D motion reconstruction. To alleviate this issue, we introduce a distillation loss that encourages the model to efficiently transfer knowledge from full-slice inputs to partial-slice inputs during training. In short, we predict two versions of motion features from full-slice observation and its randomly sampled subset. The distillation loss encourages the encoded feature from partial slices to be close to the feature from full-slice input. We formulate the loss we minimize to achieve this in Section 3.5.
@@
$V^{0}_{\rm static}=AFA(F^{0}_{2d},F^{0}_{2d})=\text{MHAttention}(Q,K,V)\;,$ (3)
@@
Note that the initial 2D encoding block and the subsequent 3D convolutions are shared for both static and motion feature extraction. This parameter sharing strategy enabled by our unified design allows motion prediction to benefit from the rich 3D spatial information learned in the static reconstruction task. Consequently, it facilitates more accurate motion estimation—an advantage overlooked by some previous works [13, 14]. As demonstrated in the experiments, this architecture leads to superior motion reconstruction performance.

### v2

# End-to-End 4D Heart Recovery Across Full-Stack and Sparse Cardiac MRI

@@ Abstract
Reconstructing cardiac motion from CMR sequences is critical for diagnosis, prognosis, and intervention. Existing methods rely on complete CMR stacks to infer full heart motion, limiting their applicability during intervention when only sparse observations are available. We present TetHeart, the first end-to-end framework for unified 4D heart mesh recovery from both offline full-stack and intra-procedural sparse-slice observations. Our method leverages deformable tetrahedra to capture shape and motion in a coherent space shared across cardiac structures. Before a procedure, it initializes detailed, patient-specific heart meshes from high-quality full stacks, which can then be updated using whatever slices can be obtained in real-time, down to a single one during the procedure. TetHeart incorporates several key innovations: (i) an attentive slice-adaptive 2D–3D feature assembly mechanism that integrates information from arbitrary numbers of slices at any position; (ii) a distillation strategy to ensure accurate reconstruction under extreme sparsity; and (iii) a weakly supervised motion learning scheme requiring annotations only at keyframes, such as the end-diastolic and end-systolic phases. Trained and validated on three large public datasets and evaluated zero-shot on additional private interventional and public datasets without retraining, TetHeart achieves state-of-the-art accuracy and strong generalization in both pre- and intra-procedural settings.
@@ 1 Introduction
Modeling cardiac shape and motion is a vital 4D reconstruction problem with direct implications for medical imaging, motion analysis, and image-guided intervention [39, 22, 33, 32]. Cardiac magnetic resonance (CMR) provides high-quality temporal anatomy and has been widely used for both shape reconstruction [2, 51, 14, 50, 48] and motion estimation [34, 26, 27, 28, 52]. However, as shown in Fig. 1 (left), existing methods rely on full volumetric CMR stacks and are thus restricted to offline analysis and pre-operative planning. In interventional settings, as shown in Fig. 1 (right), only a handful of low-resolution slices can be acquired in real time [16, 29, 36], a limitation we observed in practice at our collaborating university hospital.
This makes current 4D modeling pipelines unusable during intervetions, which motivated us to go beyond the current state of the art and to recover full 3D shape and motion from sparse and partial observations. This will enable real-time tracking of patient-specific heart dynamics, with a view to delivering guidance to interventional teams.
Figure 1: Difference between the standard offline scenario and the intervention room setting for cardiac motion reconstruction.
To this end, we introduce TetHeart, a unified framework for 4D cardiac motion reconstruction that operates consistently across both full-stack and sparse-slice CMR inputs, the latter being what can be acquired in real-time during an intervention. As shown in Fig. 2, TetHeart first constructs a patient-specific mesh representation from pre-operative data and then dynamically updates it during procedures using only the slices that can be acquired in real-time, down to a single one. TetHeart builds on deep marching tetrahedra [40], which combines the optimization efficiency of signed distance fields with the geometric flexibility of tetrahedral meshes, providing a coherent space for joint shape–motion learning. Given this, our main contributions are as follows:
  • We propose a slice-adaptive 2D–3D feature fusion mechanism that allows our networks to dynamically aggregate features from arbitrarily located slices, enabling robust reconstruction across varying input sparsity.
  • We introduce a full-to-sparse distillation strategy that transfers knowledge from dense to sparse regimes, ensuring accurate reconstruction under extreme input sparsity.
  • We develop a weakly supervised motion learning scheme that requires annotations only at key cardiac phases (end-diastolic and end-systolic) and learns motion through temporal consistency. This is essential in practice because it makes the training feasible and affordable across diverse clinical environments.
  • We demonstrate that a single model trained once generalizes seamlessly across offline and online scenarios, achieving state-of-the-art results on ACDC [5], M&Ms [9], and M&Ms-2 [25]. It also shows strong generalizability by further validating on a privately collected interventional CMR dataset and the public 4DM [52] dataset without retraining, demonstrating its potential for deployment in real-world clinical environments.
In short, the combination of our feature fusion mechanism and training scheme yield a novel unified framework that enables consistent reconstruction and motion inference across a wide spectrum of observational sparsity. As a result, TetHeart can operate both offline and online, thus enabling comprehensive, patient-specific motion tracking across the entire clinical workflow. We achieve state-of-the-art accuracy for offline use with full-stack CMR, and perhaps more importantly, we unlock online use in intra-procedural scenarios where only sparse observations are available, which cannot be handled using current methods. We will make our code publicly available.
@@
Figure 2: TetHeart reconstructs cardiac motion from observed 2D slices. At $t=0$, a 3D CMR stack $\mathbf{I}^{0}$ is processed by the AFA module with self-attention to extract a volumetric static feature $V_{\text{static}}^{0}$, which is then used to generate the initial tetrahedra $\mathcal{G}^{0}$. For each subsequent time step $t$, the observed 2D slices $\mathbf{O}^{t}$ are encoded via the AFA module with cross-attention to obtain a volumetric motion feature $V_{\text{motion}}^{t}$, which is then used to recover motion by predicting offsets that deform $\mathcal{G}^{0}$. Weights are shared between the static and dynamic branch.
@@ 2 Related Work
We first discuss current approaches to recovering static and dynamic hearts, and then the imaging modality we focus on.
@@ 2.1 Static Cardiac Shape Modeling
Reconstructing the heart from 3D medical images holds significant clinical value. However, CMR images often exhibit low through-plane resolution, making accurate reconstruction challenging. Current solutions [42, 2, 45, 15, 1, 10, 47, 4, 18, 48] use either traditional optimization-based techniques or deep learning-based methods.
Optimization-Based Approaches usually rely on a two-stage pipeline, often starting from contours manually extracted from the CMR images. For example, [45] starts from a predefined tubular mesh that is deformed to minimize a contour-matching loss and, in [18], a 3D active shape model and an intensity model are used to align initial meshes with GT contours. Even though this works, relying on test-time optimization makes the reconstruction time-consuming, taking from tens of seconds to several minutes per frame. When applied to a full image sequence, this may extend to over an hour, making it impractical for real-time applications. Moreover, it has only been evaluated using GT contours, requiring labor-intensive manual annotation.
Deep Learning-Based Approaches leverage prior knowledge learned from large datasets to accelerate mesh reconstruction. For instance, in [4], contours are treated as sparse point cloud and decoded into 3D meshes using point- and graph-convolutions. Similarly, in [10], contours are embedded into a 3D volume and a model made of 3D CNNs and GCNs is used to progressively deform a template mesh. These methods are much faster but also have drawbacks. Reconstructing meshes solely from extracted contours or points fails to exploit the rich appearance information contained in cine images [48]. Furthermore, as contour extraction is not necessarily perfect, such two-step approaches are subject to error accumulation [10].
@@ 2.2 Dynamic Cardiac Motion Modeling
Arguably, when reconstructing deforming hearts from sequences of CMR scans, one could simply run a static reconstruction algorithm on each individual frame. However, this would fail to exploit temporal consistency and make it difficult to establish inter-frame correspondences. Thus, several motion-based approaches that compute the deformation from frame to frame have been proposed.
@@
When full CMR stacks are available, recent deep learning-based approaches [26, 27, 28, 52, 51] can be used. MulViMotion [27] trains 3D CNNs to estimate left ventricular myocardial motion by predicting voxel-based deformation fields from CMR data, and then building 3D meshes by warping segmentation masks. DeepMesh [28] decouples mesh reconstruction from motion estimation. It uses 3D CNNs and interpolation to predict per-vertex deformation. 4DMR [52] models the myocardium using implicit representation. Decoupled shape and motion latent codes are optimized at test time to predict a source shape and motion fields. We use MulViMotion, DeepMesh and 4DMR as baselines because they are among the best current methods.
@@
Unfortunately, these approaches also suffer from several limitations. First, some of them focus solely on the myocardium, neglecting other cardiac structures and preventing comprehensive heart assessment. Second, optimization-based methods like 4DMR are time-consuming at inference time. Finally and most damagingly, they require full scans that cannot be acquired at sufficient frame rate, which precludes intra-procedural use as discussed below.
2.3 Real-Time MR Imaging
@@
MRI has a reputation for being slow. Acquiring a full MRI stack involves sequential slice acquisition over a 4–6 heartbeat breath-hold, which can take several minutes [29, 36]. While this is feasible for pre-interventional planning, but impractical for intra-procedural work.
@@
Recently, real-time MRI (RT-MRI) has gained favor due to its ability to capture dynamic processes without gating, synchronization, or repetition [29]. Its applications include complex electrophysiology procedures [30], congenital heart disease [17], and cardiac interventions [16, 49]. RT-MRI provides excellent soft-tissue contrast for assessing interventions like ablations [6, 36] and allows for continuous visualization and tracking of devices, and it is free of ionizing radiation [30, 8].
@@
Although RT-MRI enables rapid continuous image acquisition, it still involves a fundamental tradeoff between spatial resolution, temporal resolution, artifacts and reconstruction latency [29]. Therefore, achieving a sufficient temporal resolution (e.g., 20–25 Hz) for use during invasive procedures, can only be done at the cost of acquiring a limited number of 2D slices at lower spatial resolution [16, 49, 8, 29, 36]. We are not aware of any current method that can reconstruct 3D cardiac motion using this kind of data. Yet, delivering a high-quality, dynamic 3D heart model to the intervention room based on real-time observations is crucial for guiding high-precision minimally invasive cardiac procedures [23, 19, 16] and for real-time localization of devices within the heart [30, 41]. This is the application that motivated us to begin this study.
@@ 3 Method
We now present TetHeart, a framework for accurate 4D cardiac shape and motion reconstruction from either full CMR volumes or sparse 2D slice sets, as shown in Fig. 2. We first describe our heart representation based on deep marching tetrahedra [40], an explicit–implicit hybrid model combining the optimization efficiency of signed distance fields with the geometric flexibility of tetrahedral meshes. Next, we introduce the Attentive 2D–3D Feature Assembler (AFA) we designed to adaptively aggregate features from arbitrary numbers of 2D slices at varying spatial locations. Finally, we detail our weakly supervised training scheme, enabling effective motion learning from limited annotations available only at key cardiac phases.
@@ 3.1 Formalization
To model deforming hearts, we use the deep marching tetrahedra formalism. We discretize the 3D space using a deformable tetrahedral grid, in which each vertex possesses an SDF, denoted as $\mathcal{G}=(\{\mathbf{v}_{i}\},\{s_{i}\},\mathcal{T})$, where $\mathbf{v}_{i}$ and $s_{i}$ are the vertices and their corresponding SDFs in $\mathcal{T}$, the set of all tetrahedrons. The explicit surface can then be extracted using the differentiable marching tetrahedra algorithm [13].
Given an initial 3D heart model $\mathcal{G}^{0}$ reconstructed from the end-diastolic (ED) frame $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$, our goal is to estimate a deformed model $\mathcal{G}^{t}$ at each time instant $t$, based on subsequent observations $\mathbf{O}^{t}$ that reflect the ongoing cardiac motion. We do this by updating the vertices of $\mathcal{G}^{0}$:
@@
    Online. For intra-procedural use, $\mathbf{I}^{0}\in\mathbb{R}^{D\times H\times W}$ is a full-slice image acquired before the intervention, taking as much time as needed. However, it has to be possible to acquire $\mathbf{O}^{t}$ at sufficient frame rate and, consequently, it can only comprise a small set—from 1 to 3—of 2D slices at lower spatial resolution and at arbitrary locations [16].
@@ 3.2 Attentive 2D-3D Feature Assembler
Because the observations can be either a full stack, a sparse set of slices, or anything in between, we need a mechanism to extract useful features in all cases. To this end, we developed the Attentive 2D-3D Feature Assembler (AFA) module that takes $\mathbf{I}^{0}$ and $\mathbf{O}^{t}$ as input and produces features.
Let us write $\mathbf{O}^{t}$ as $\{\mathbf{I}^{t}_{s}\in\mathbb{R}^{H\times W}|s=1,\ldots,S\}$, a collection of 2D slices where $S$ is the number of slices. With this, $\mathbf{I}^{0}$ can be written as $\{\mathbf{I}^{0}_{d}\in\mathbb{R}^{H\times W}|d=1,\ldots,D\}$. A full-slice observation is simply one where $S=D$.
@@
Motion Features. A naive way to use attention to construct motion features would be to treat the $F^{0}_{2d}$ as the query and the $F^{t}_{2d}$ as both keys and values. However, this would be inefficient because global attention aims at capturing long-range dependencies, which is not needed to model the inherently local nature of anatomical motion. Also, this requires computing full pairwise relationships between all query and key positions, which is computationally expensive when using many slices and high-resolution feature maps.
To avoid this, we take our inspiration from the Swin Transformer [24] and use instead a variant of attention that preserves locality and significantly reduces computational complexity. For each voxel in $F^{0}_{2d}$, we first identify $k_{1}$ spatially nearest 2D slices, or all slices if there are only $k_{1}$ or less. Within each selected slice, we then retrieve the features at the $k_{2}$ spatially closest positions to serve as keys and values. Attention is subsequently computed only between the query and its corresponding set of localized keys and values. This can be written as
$V^{t}_{\rm motion}=AFA(F^{0}_{2d},F^{t}_{2d})=\text{MHAttention}(Q,K,V)\;,$
@@
where NN-Select denotes the nearest slice and position selecting operation. MHAttention denotes a Multi-Head Attention [43] operation. A learnable position embedding is used to encode the spatial information. In practice, we set $k_{1}=3$ and $k_{2}=9$. This configuration ensures that the computational complexity of our attention mechanism is approximately equivalent to that of a 3D convolution with a kernel size of 3. As a result, the AFA module achieves high computational efficiency while retaining the flexibility and adaptability of attention-based feature aggregation.
While our AFA module effectively encodes features from varying 2D slice configurations, a persistent challenge remains in the online scenario, where the limited number of slices can hinder accurate 3D motion reconstruction. To alleviate this issue, we introduce a distillation loss that encourages the model to transfer knowledge from full-slice inputs to partial-slice inputs during training. In short, we predict two versions of motion features from full-slice observation and its randomly sampled subset. The distillation loss encourages the encoded feature from partial slices to be close to the feature from full-slice input. We formulate the loss we minimize to achieve this in Section 3.5.
@@
$V^{0}_{\rm static}=AFA(F^{0}_{2d},F^{0}_{2d})=\text{MHAttention}(Q,K,V)\;,$
@@
Note that the initial 2D encoding block and subsequent 3D convolutions are shared for static and motion feature extraction. This unified design allows motion prediction to benefit from the rich 3D spatial information learned in the static reconstruction task, facilitating more accurate motion estimation—an advantage overlooked by some previous works [28, 52]. As shown in the experiments, this design leads to superior motion reconstruction performance.